# AI and the Net-Zero Journey: Energy Demand, Emissions, and the Potential for Transition

### v1

@@ 5 The Energy Consumption Scenarios and Near-Term Outlook for Data Centers and AI
    Horizon - High Efficiency Case - Sustainable AI: The Horizon scenario focuses on achieving net-zero emissions by 2050 and limiting global warming to 1.5°C, emphasizing sustainability and efficiency in energy use. Similarly, the High Efficiency Case from the IEA assumes significant improvements in energy efficiency, both in AI technologies and the broader energy sector, envisioning reduced energy consumption per unit of AI output. The Sustainable AI scenario by Paccou and Wijnhoven (2024) balances AI development with sustainability, highlighting the importance of generative AI inferencing and traditional AI in decarbonization efforts. These scenarios share an emphasis on achieving sustainability and efficiency in AI and energy consumption. Under these scenarios, energy demand by data centers approaching a plateau by 2035 at nearly 1000 TWh. AI accounts for approximately 80% of the electricity usage (Figure 6c).
@@
    Baseline Case - Energy Crisis: The Baseline Case from the IEA assumes current trends and policies continue without significant changes, serving as a benchmark for comparing other scenarios and representing a future where AI development follows existing trajectories. The Energy Crisis scenario by Paccou and Wijnhoven (2024) considers the possibility of an energy crisis caused by AI, highlighting the risks of unmanaged AI development leading to significant energy consumption and potential shortages. It emphasizes the need for strategic planning and management to avoid such crises. These scenarios share a focus on the potential risks and challenges associated with AI development and energy consumption. Under these scenarios, energy demand by data centers approaching a plateau in 2035 at around 1200 TWh. AI electricity usage reaches its maximum in 2030 where it accounts for 70% of the data center’s electricity usage. However, due to the energy crisis, it drops to around 16% by 2035 (Figure 6d).
@@ 6 The Near-Term Carbon Footprint: AI as a Net-Contributor
Although energy consumption scenarios for data centers are based on global estimates, they can greatly vary by country. This leads to a diverse global landscape influenced by global trends, local conditions, policies, and decisions. In fact, geopolitical factors play a crucial role in shaping AI’s impact on the energy transition and carbon footprint in the short term. Regulatory frameworks across different countries can either accelerate or hinder AI advancements, influencing how quickly AI technologies are adopted in energy sectors. National security concerns drive investments in AI for defense and surveillance, which can divert resources from sustainable energy initiatives. International trade policies and technological alliances also affect the global distribution of AI technologies, impacting their deployment in renewable energy projects [wef (2025)]. Control over semiconductor supply chains, essential for AI hardware, can create geopolitical tensions, affecting the availability and cost of AI technologies. Additionally, countries leading in AI development, such as the US and China, have significant influence over global AI policies and standards, shaping the direction of AI applications in energy efficiency and carbon reduction. These geopolitical dynamics, combined with societal factors like public perception and workforce adaptation, will determine AI’s role in achieving a sustainable energy transition and reducing carbon emissions [Iqbal et al. (2025)].
Despite the growing emphasis on renewable energy, many regions still rely heavily on fossil fuels to power their electricity grids. This means that a significant portion of data centers’ energy consumption is indirectly contributing to GHG emissions. In regions with limited access to clean energy sources, the increasing electricity demands of data centers can exacerbate the reliance on fossil fuels, hindering efforts to decarbonize the energy sector [Maher et al. (2022)].
@@
More precisely, the CO2 emissions by data centers in the Headwinds Case and the High Efficiency Case reached plateau by 2030 at around 130-150 Mt CO2 per year. In the High Efficiency Case the trend started pointing downwards by 2035. Finally, in the Base and Lift-Off Cases, the CO2 emissions continued to increase past 2035 albeit at different rates reaching 210 and 300 Mt per year respectively (Figure 7a). In the case of AI, the CO2 emissions from electricity usage in Sustainable AI and Limits to Growth plateaued around 2030 at around 115 Mt per year. On the contrary, the Abundance without Boundaries scenario grew at a much faster pace reaching 240 Mt CO2 by 2035 with no clear signs of slowing down. Finally, the Energy Crisis scenario peaked at 130 Mt CO2 in 2030 followed by a rapid decline to 35 Mt within five years (Figure 7b). For context, 100 Mt CO2 is equivalent to the emissions from nearly 13.5 million American home’s energy usage for one year (epa (2025)).
Figure 7: Historical trends and forecasts for CO2 emissions from a) data centers and b) AI workloads through 2035. These forecasts are based on the data provided in the She (2025), 2024 (2025) and Paccou and Wijnhoven (2024)
As illustrated in Figure 7, these scenarios present contrasting perspectives on the penetration of AI and its impact on CO2 emissions. For example, the Lift-Off Case from IEA is more optimistic, projecting the highest CO2 emissions, whereas the Energy Crisis scenario from Paccou and Wijnhoven (2024) is pessimistic, predicting the least emissions. Although these reports present extreme scenarios, we believe the reality lies somewhere in between. It is often preferable to provide an interval within which the forecasted CO2 emissions are expected to fall. To determine this interval, we consider the mean scenario presented in Figure 7 and calculate the P5, P50, and P95 probability cases around that, which will be described next.
@@
Figure 8: The P5, P50, P95 probability scenarios generated based on the following assumptions: a) The input variables listed in Table 2 were assumed to follow a normal distribution. The ranges for these variables were selected within a 99% confidence interval around the mean scenario. b) The scenarios presented in Paccou and Wijnhoven (2024) were reproduced without considering the energy crisis scenario.
@@ 7 The Long-Term Promise: AI as a Net-Zero Enabler
The increasing presence of AI raises concerns about its energy consumption. However, it also has the potential to reduce emissions and help achieve climate targets. AI applications can optimize energy usage in sectors such as transportation, manufacturing, and agriculture. For example, AI-powered smart grids can enhance grid efficiency, minimize energy loss during transmission, and seamlessly integrate renewable energy sources. Studies indicate that intelligent-applications coupled with smart grids could reduce energy consumption by 15% [Jin et al. (2024)]. AI-based applications are already enabling predictive maintenance, anticipating equipment failures, thereby reducing downtime and energy waste. Furthermore, AI can optimize supply chains, leading to decreased transportation emissions. AI can accelerate project delivery timelines for CCS through very fast surrogate-based modeling. It is difficult to gauge the overall impact of AI on CO2 reductions as there are many technological, socio-economic, and political factors influencing that journey. A recent study from Google and BCG suggests that implementing AI for corporate sustainability could result in a substantial 2.6 to 5.3 gigaton reduction in CO2 equivalent by 2030 if potential optimizations were to fully penetrate industry [Maher et al. (2022)]. Similarly, [2024 (2025)] reported that AI applications in the energy sector could lead to a reduction of 1.4 gigaton of CO2 by 2035 in the Widespread Adoption Case.
@@
Nuclear Power: AI is aiding nuclear energy in reducing emissions through various applications. AI is being used in optimizing the designs and operation of small modular reactors (SMRs) and their integration into the energy grid. AI algorithms are finding ways to enhance the efficiency of SMRs by predicting the most effective operational parameters, thus maximizing energy output while minimizing waste. By optimizing reactor operations, AI enhances energy generation efficiency [Alatalo et al. (2025)]. Predictive maintenance powered by AI helps foresee equipment failures, reducing downtime and ensuring peak operational efficiency [Myers et al. (2025)]. Accurate energy demand forecasting allows for better planning, reducing reliance on backup fossil fuel plants [Alatalo et al. (2025)]. AI also enhances safety by monitoring reactor data in real-time to identify anomalies, thus preventing potential accidents [Myers et al. (2025)]. Additionally, AI improves grid management, ensuring efficient energy distribution and lowering fossil fuel dependence [Khym and Vasquez (2025)].
@@
Wind and Solar: AI is playing a transformative role in reducing emissions from wind and solar energy by optimizing various aspects of these renewable energy systems. In wind energy, AI enhances efficiency through precise weather forecasting and wind analysis, allowing for better planning and operational adjustments. AI-driven predictive maintenance helps identify potential issues before they become significant problems, reducing downtime and ensuring turbines operate at peak efficiency. Additionally, AI-powered tools for turbine monitoring and inspection improve safety and performance by detecting defects that might be missed by human inspectors [Dörterler et al. (2024), Morkos (2025)]. In solar energy, AI maximizes energy output by analyzing data from various sources, such as weather conditions and system performance, to predict energy production and optimize energy storage. Predictive analytics and machine learning algorithms enhance the efficiency of solar panels by managing factors like shading and angle adjustments [Gul et al. (2025), SolarAI (2025)]. AI also plays a crucial role in improving grid management, ensuring that the energy generated from wind and solar sources is distributed efficiently, reducing reliance on fossil fuels and lowering overall emissions [Khym and Vasquez (2025)].
@@
Biofuels: By leveraging AI and machine learning algorithms, researchers can optimize the entire biofuel production process, from feedstock selection to final fuel production. AI helps in analyzing large datasets to identify the most efficient feedstocks and processing conditions, thereby maximizing yield and minimizing environmental impact. Predictive models developed using AI can forecast the performance of different biofuel production methods, allowing for better planning and resource allocation [Okolie (2024)]. Additionally, AI-driven automation in bioprocessing can enhance the efficiency and consistency of biofuel production, reducing waste and emissions. AI also aids in the development of advanced catalysts and enzymes that can accelerate biofuel production while lowering energy consumption [2023 (2023), Salama (2025)].
@@
Hydrogen: Green hydrogen energy, produced through water electrolysis, has emerged as a crucial technology in this transition. Proton exchange membrane water electrolyzers (PEMWEs) have gained significant attention due to their efficiency and compatibility with renewable energy sources, positioning them as a frontrunner in green hydrogen production technologies. Challenges with PEMWEs include the need for innovative research to overcome existing limitations in key materials such as electrocatalysts, membranes, membrane electrode assemblies (MEAs), stacks, and systems. AI’s transformative potential in PEMWE research is focusing on discovery and optimization of key materials, as well as the design and management of MEAs, cells, and stacks [Wang et al. (2025)]. AI-enabled simulations provide valuable insights during feasibility studies, helping to determine the optimal number of electrolyzers, capacity utilization, and energy demand for hydrogen production. These simulations also incorporate renewable energy models, weather forecasts, and electricity market prices to optimize the entire lifecycle of hydrogen projects. Furthermore, AI plays a vital role in managing energy grids, ensuring efficient integration of hydrogen energy and reducing reliance on fossil fuels [Mullanu et al. (2024)].
@@
Carbon Capture and Storage (CCS): Several key technological advancements in CCS technologies that are crucial for reducing the carbon footprint of energy-intensive industries, including AI. One significant advancement is the development of smaller, more efficient capture systems. These systems are designed to be more compact and cost-effective, making them suitable for a wider range of applications. For example, the use of modular capture units that can be easily integrated into existing industrial processes. These units are capable of capturing CO2 emissions at a lower cost and with higher efficiency compared to traditional systems. Another important development is the integration of CCS with renewable energy sources. This approach involves using renewable energy to power the capture and storage processes, thereby reducing the overall carbon footprint of the CCS operations. For instance, some CCS projects are now being powered by solar or wind energy, which not only reduces emissions but also enhances the sustainability of the entire process. New storage techniques, such as mineral carbonation and enhanced oil recovery (EOR), are being developed with the help of AI to improve the efficiency and safety of CO2 storage. Recent advances in operator learning are accelerating flow simulations conducted to quantify the CO2 plume migration by three to four orders of magnitude [Pawar et al. (2025), Chandra et al. (2024)]. These technological advancements are crucial for reducing the carbon footprint of energy-intensive industries, including AI. By making CCS more efficient, cost-effective, and sustainable, these innovations help mitigate the environmental impact of AI and other advanced technologies, contributing to global climate goals [2024 (2024)].
@@
As part of the energy system decarbonization, AI can also transform transportation, a major source of emissions. Self-driving vehicles, optimized traffic systems, and improved logistics planning have the potential to significantly reduce our reliance on fossil fuels. By analyzing real-time data, AI can forecast traffic trends, streamline routes, and reduce congestion. Research on traffic prediction encompasses various models with examples including feature-based models [Li et al. (2019)], Gaussian process models [Salinas et al. (2019)], state-space models [Duan et al. (2019)], and deep learning models [Yin et al. (2022)]. Autonomous vehicle development is another nascent area for AI [Bojarski et al. (2016),Wu et al. (2017)] predominantly targeting safety and energy efficiency. AI-driven smart mobility systems has the potential to improve the energy efficiency of public transportation through dynamic scheduling to guarantee timely arrivals and departures. Finally. AI can accelerate the discovery of sustainable materials for vehicles and associated manufacturing processes. Another important element of the energy system is heating and cooling of buildings. AI models can process sensor data at the building and city level to optimize heating and cooling systems, lighting, and energy consumption patterns [Qolomany et al. (2019), Pham et al. (2020), Zekić-Sušac et al. (2021), Alshurideh et al. (2023)] leading to energy savings and reduction in carbon emissions [Milojevic-Dupont and Creutzig (2021), Zhang et al. (2023)].

### v2

@@ 5 The Energy Consumption Scenarios and Near-Term Outlook for Data Centers and AI
    Horizon - High Efficiency Case - Sustainable AI: The Horizon scenario focuses on achieving net-zero emissions by 2050 and limiting global warming to 1.5°C, emphasizing sustainability and efficiency in energy use. Similarly, the High Efficiency Case from the IEA assumes significant improvements in energy efficiency, both in AI technologies and the broader energy sector, envisioning reduced energy consumption per unit of AI output. The Sustainable AI scenario by paccouandwijnhoven2024 balances AI development with sustainability, highlighting the importance of generative AI inferencing and traditional AI in decarbonization efforts. These scenarios share an emphasis on achieving sustainability and efficiency in AI and energy consumption. Under these scenarios, energy demand by data centers approaching a plateau by 2035 at nearly 1000 TWh. AI accounts for approximately 80% of the electricity usage (Figure 6c).
@@
    Baseline Case - Energy Crisis: The Baseline Case from the IEA assumes current trends and policies continue without significant changes, serving as a benchmark for comparing other scenarios and representing a future where AI development follows existing trajectories. The Energy Crisis scenario by paccouandwijnhoven2024 considers the possibility of an energy crisis caused by AI, highlighting the risks of unmanaged AI development leading to significant energy consumption and potential shortages. It emphasizes the need for strategic planning and management to avoid such crises. These scenarios share a focus on the potential risks and challenges associated with AI development and energy consumption. Under these scenarios, energy demand by data centers approaching a plateau in 2035 at around 1200 TWh. AI electricity usage reaches its maximum in 2030 where it accounts for 70% of the data center’s electricity usage. However, due to the energy crisis, it drops to around 16% by 2035 (Figure 6d).
@@ 6 The Near-Term Carbon Footprint: AI as a Net-Contributor
Although energy consumption scenarios for data centers are based on global estimates, they can greatly vary by country. This leads to a diverse global landscape influenced by global trends, local conditions, policies, and decisions. In fact, geopolitical factors play a crucial role in shaping AI’s impact on the energy transition and carbon footprint in the short term. Regulatory frameworks across different countries can either accelerate or hinder AI advancements, influencing how quickly AI technologies are adopted in energy sectors. National security concerns drive investments in AI for defense and surveillance, which can divert resources from sustainable energy initiatives. International trade policies and technological alliances also affect the global distribution of AI technologies, impacting their deployment in renewable energy projects [weforumpodcast2024]. Control over semiconductor supply chains, essential for AI hardware, can create geopolitical tensions, affecting the availability and cost of AI technologies. Additionally, countries leading in AI development, such as the US and China, have significant influence over global AI policies and standards, shaping the direction of AI applications in energy efficiency and carbon reduction. These geopolitical dynamics, combined with societal factors like public perception and workforce adaptation, will determine AI’s role in achieving a sustainable energy transition and reducing carbon emissions [Iqbal2025].
Despite the growing emphasis on renewable energy, many regions still rely heavily on fossil fuels to power their electricity grids. This means that a significant portion of data centers’ energy consumption is indirectly contributing to GHG emissions. In regions with limited access to clean energy sources, the increasing electricity demands of data centers can exacerbate the reliance on fossil fuels, hindering efforts to decarbonize the energy sector [bcg_climate_2022].
@@
More precisely, the CO2 emissions by data centers in the Headwinds Case and the High Efficiency Case reached plateau by 2030 at around 130-150 Mt CO2 per year. In the High Efficiency Case the trend started pointing downwards by 2035. Finally, in the Base and Lift-Off Cases, the CO2 emissions continued to increase past 2035 albeit at different rates reaching 210 and 300 Mt per year respectively (Figure 7a). In the case of AI, the CO2 emissions from electricity usage in Sustainable AI and Limits to Growth plateaued around 2030 at around 115 Mt per year. On the contrary, the Abundance without Boundaries scenario grew at a much faster pace reaching 240 Mt CO2 by 2035 with no clear signs of slowing down. Finally, the Energy Crisis scenario peaked at 130 Mt CO2 in 2030 followed by a rapid decline to 35 Mt within five years (Figure 7b). For context, 100 Mt CO2 is equivalent to the emissions from nearly 13.5 million American home’s energy usage for one year (epacalculator).
Figure 7: Historical trends and forecasts for CO2 emissions from a) data centers and b) AI workloads through 2035. These forecasts are based on the data provided in the Shell_Sce_2025, iea2025 and paccouandwijnhoven2024
As illustrated in Figure 7, these scenarios present contrasting perspectives on the penetration of AI and its impact on CO2 emissions. For example, the Lift-Off Case from IEA is more optimistic, projecting the highest CO2 emissions, whereas the Energy Crisis scenario from paccouandwijnhoven2024 is pessimistic, predicting the least emissions. Although these reports present extreme scenarios, we believe the reality lies somewhere in between. It is often preferable to provide an interval within which the forecasted CO2 emissions are expected to fall. To determine this interval, we consider the mean scenario presented in Figure 7 and calculate the P5, P50, and P95 probability cases around that, which will be described next.
@@
Figure 8: The P5, P50, P95 probability scenarios generated based on the following assumptions: a) The input variables listed in Table 2 were assumed to follow a normal distribution. The ranges for these variables were selected within a 99% confidence interval around the mean scenario. b) The scenarios presented in paccouandwijnhoven2024 were reproduced without considering the energy crisis scenario.
@@ 7 The Long-Term Promise: AI as a Net-Zero Enabler
The increasing presence of AI raises concerns about its energy consumption. However, it also has the potential to reduce emissions and help achieve climate targets. AI applications can optimize energy usage in sectors such as transportation, manufacturing, and agriculture. For example, AI-powered smart grids can enhance grid efficiency, minimize energy loss during transmission, and seamlessly integrate renewable energy sources. Studies indicate that intelligent-applications coupled with smart grids could reduce energy consumption by 15% [Jin2024]. AI-based applications are already enabling predictive maintenance, anticipating equipment failures, thereby reducing downtime and energy waste. Furthermore, AI can optimize supply chains, leading to decreased transportation emissions. AI can accelerate project delivery timelines for CCS through very fast surrogate-based modeling. It is difficult to gauge the overall impact of AI on CO2 reductions as there are many technological, socio-economic, and political factors influencing that journey. A recent study from Google and BCG suggests that implementing AI for corporate sustainability could result in a substantial 2.6 to 5.3 gigaton reduction in CO2 equivalent by 2030 if potential optimizations were to fully penetrate industry [bcg_climate_2022]. Similarly, [iea2025] reported that AI applications in the energy sector could lead to a reduction of 1.4 gigaton of CO2 by 2035 in the Widespread Adoption Case.
@@
Nuclear Power: AI is aiding nuclear energy in reducing emissions through various applications. AI is being used in optimizing the designs and operation of small modular reactors (SMRs) and their integration into the energy grid. AI algorithms are finding ways to enhance the efficiency of SMRs by predicting the most effective operational parameters, thus maximizing energy output while minimizing waste. By optimizing reactor operations, AI enhances energy generation efficiency [app15020999]. Predictive maintenance powered by AI helps foresee equipment failures, reducing downtime and ensuring peak operational efficiency [MYERS2025105568]. Accurate energy demand forecasting allows for better planning, reducing reliance on backup fossil fuel plants [app15020999]. AI also enhances safety by monitoring reactor data in real-time to identify anomalies, thus preventing potential accidents [MYERS2025105568]. Additionally, AI improves grid management, ensuring efficient energy distribution and lowering fossil fuel dependence [yale2025].
@@
Wind and Solar: AI is playing a transformative role in reducing emissions from wind and solar energy by optimizing various aspects of these renewable energy systems. In wind energy, AI enhances efficiency through precise weather forecasting and wind analysis, allowing for better planning and operational adjustments. AI-driven predictive maintenance helps identify potential issues before they become significant problems, reducing downtime and ensuring turbines operate at peak efficiency. Additionally, AI-powered tools for turbine monitoring and inspection improve safety and performance by detecting defects that might be missed by human inspectors [dorterler2024, morkos2025]. In solar energy, AI maximizes energy output by analyzing data from various sources, such as weather conditions and system performance, to predict energy production and optimize energy storage. Predictive analytics and machine learning algorithms enhance the efficiency of solar panels by managing factors like shading and angle adjustments [GUL2025125210, solarai]. AI also plays a crucial role in improving grid management, ensuring that the energy generated from wind and solar sources is distributed efficiently, reducing reliance on fossil fuels and lowering overall emissions [yale2025].
@@
Biofuels: By leveraging AI and machine learning algorithms, researchers can optimize the entire biofuel production process, from feedstock selection to final fuel production. AI helps in analyzing large datasets to identify the most efficient feedstocks and processing conditions, thereby maximizing yield and minimizing environmental impact. Predictive models developed using AI can forecast the performance of different biofuel production methods, allowing for better planning and resource allocation [OKOLIE2024100928]. Additionally, AI-driven automation in bioprocessing can enhance the efficiency and consistency of biofuel production, reducing waste and emissions. AI also aids in the development of advanced catalysts and enzymes that can accelerate biofuel production while lowering energy consumption [doe2023, Salama2025].
@@
Hydrogen: Green hydrogen energy, produced through water electrolysis, has emerged as a crucial technology in this transition. Proton exchange membrane water electrolyzers (PEMWEs) have gained significant attention due to their efficiency and compatibility with renewable energy sources, positioning them as a frontrunner in green hydrogen production technologies. Challenges with PEMWEs include the need for innovative research to overcome existing limitations in key materials such as electrocatalysts, membranes, membrane electrode assemblies (MEAs), stacks, and systems. AI’s transformative potential in PEMWE research is focusing on discovery and optimization of key materials, as well as the design and management of MEAs, cells, and stacks [WANG2025]. AI-enabled simulations provide valuable insights during feasibility studies, helping to determine the optimal number of electrolyzers, capacity utilization, and energy demand for hydrogen production. These simulations also incorporate renewable energy models, weather forecasts, and electricity market prices to optimize the entire lifecycle of hydrogen projects. Furthermore, AI plays a vital role in managing energy grids, ensuring efficient integration of hydrogen energy and reducing reliance on fossil fuels [MULLANU2024].
@@
Carbon Capture and Storage (CCS): Several key technological advancements in CCS technologies that are crucial for reducing the carbon footprint of energy-intensive industries, including AI. One significant advancement is the development of smaller, more efficient capture systems. These systems are designed to be more compact and cost-effective, making them suitable for a wider range of applications. For example, the use of modular capture units that can be easily integrated into existing industrial processes. These units are capable of capturing CO2 emissions at a lower cost and with higher efficiency compared to traditional systems. Another important development is the integration of CCS with renewable energy sources. This approach involves using renewable energy to power the capture and storage processes, thereby reducing the overall carbon footprint of the CCS operations. For instance, some CCS projects are now being powered by solar or wind energy, which not only reduces emissions but also enhances the sustainability of the entire process. New storage techniques, such as mineral carbonation and enhanced oil recovery (EOR), are being developed with the help of AI to improve the efficiency and safety of CO2 storage. Recent advances in operator learning are accelerating flow simulations conducted to quantify the CO2 plume migration by three to four orders of magnitude [Pawar2025AcceleratedCS, chandra2024fourier]. These technological advancements are crucial for reducing the carbon footprint of energy-intensive industries, including AI. By making CCS more efficient, cost-effective, and sustainable, these innovations help mitigate the environmental impact of AI and other advanced technologies, contributing to global climate goals [gsccs2024].
@@
As part of the energy system decarbonization, AI can also transform transportation, a major source of emissions. Self-driving vehicles, optimized traffic systems, and improved logistics planning have the potential to significantly reduce our reliance on fossil fuels. By analyzing real-time data, AI can forecast traffic trends, streamline routes, and reduce congestion. Research on traffic prediction encompasses various models with examples including feature-based models [Li2019], Gaussian process models [Salinas2019], state-space models [Duan2019], and deep learning models [Yin2022]. Autonomous vehicle development is another nascent area for AI [bojarski2016,pmlr-v78-wu17a] predominantly targeting safety and energy efficiency. AI-driven smart mobility systems has the potential to improve the energy efficiency of public transportation through dynamic scheduling to guarantee timely arrivals and departures. Finally. AI can accelerate the discovery of sustainable materials for vehicles and associated manufacturing processes. Another important element of the energy system is heating and cooling of buildings. AI models can process sensor data at the building and city level to optimize heating and cooling systems, lighting, and energy consumption patterns [Qolomany2019, PHAM2020121082, ZEKICSUSAC2021102074, Ghazal2023] leading to energy savings and reduction in carbon emissions [MILOJEVICDUPONT2021102526, ZHANG2023101347].